A coil has a resistance of 50 ohms at 0celcius if the temperature coefficient of resistance for the coil is 0.0043ohms/ohms degrees celcuis determine the resistance of the coil at 25degree celcuis


1
Expert's answer
2022-05-19T05:43:11-0400
"R=R_0(1+\alpha\Delta T)\\
=50(1+0.0043*25)=55\:\Omega"